Circular No. 6426

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 K. Dennerl and J. Greiner, Max-Planck-Institut für Extraterrestrische

Physik, Garching, communicate: “The recent x-ray transient GRS 1739– 278 (IAUC 6350) was observed for 620 s with the ROSAT High Resolution Imager on Mar. 31.142 UT. A bright x-ray source with 20 counts/s was found at α = 17^{h}42^{m}40^{s}.3, δ = −27^{o}44′54′′ (equinox 2000.0; uncertainty ±8′′), consistent with the TTM position (uncertainty ±1′). The variable radio source (Hjellming et al., IAUC 6383) is inside the ROSAT error radius, suggesting it to be the radio counterpart. Although there is a faint star within the x-ray error circle, it is 6′′ off the radio position. Deep imaging is encouraged to identify the optical/infrared counterpart.”

4U 0614+091 E. Ford, P. Kaaret, and M. Tavani, Columbia University; B. A. Har-

mon, Marshall Space Flight Center (MSFC), NASA; S. N. Zhang, MSFC and Universities Space Research Association; and D. Barret, P. Bloser, and J. Grindlay, Harvard-Smithsonian Center for Astrophysics, report the detection of high-frequency, quasiperiodic oscillations (QPOs) from the xray burster 4U 0614+091 with the Rossi X-ray Timing Explorer: “Using 125-μs-resolution PCA data, two peaks are apparent in the power-density spectrum with frequencies near 900 and 570 Hz. The frequency of each peak is approximately linearly proportional to the total countrate, varying from 750 to 950 Hz (high-frequency peak) and 400 to 600 Hz (low-frequency peak) as the countrate varies from 750 to 850 counts/s (full PCA). The frequency difference between the two QPOs is constant at 327 ± 4 Hz. The relative widths (FWHM/ν) are constant at 15 and 6 percent, respectively. From gaussian fits, the amplitude above the poisson noise appears constant. A feature at 326 ± 8 Hz with marginal significance is present during one 30-min interval. This feature, if confirmed, may be direct evidence for a 3.0-ms spin period of the neutron star.”

CH CYGNI J. A. Mattei, AAVSO, reports that this symbiotic star has faded and

is now fainter than at any time on record (since 1929), as indicated by the following AAVSO observations: May 23.1 UT, 9.4 (J. Bortle, Stormville, NY); June 6.15, 9.9 (Bortle); 15.9, 10.0 (R. Fidrich, Rakoczi, Hungary); 20.30, 10.4: (G. Hanson, Cave Creek, AZ); 25.23, 10.3: (Hanson).

1996 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green